The yearbook club had a meeting. The meeting had 24 people, which is three fourths of the club. How many people are in the club?
Answer by guitarheroguy(26)

Well,
If 3/4=24, we need to find out what the other fourth is. To do this, we need to divide the total number of people by the numerator
24/8=8
Then, we add 24 and 8
24+8=32
So,
There are 32 people in the club.